# From planar to annular to toroidal bracket polynomials for pseudo knots and links

### v1

@@ Abstract.
Pseudo links are equivalence classes under Reidemeister-type moves of link diagrams containing crossings with undefined over and under information. In this paper, we extend the Kauffman bracket and Jones polynomials from planar pseudo links to pseudo links in the annulus and in the torus, and their respective lifts from the three-space to the solid torus and the thickened torus. Moreover, since annular and toroidal pseudo links can be represented as mixed links in the three-sphere, we also introduce the respective Kauffman bracket and Jones polynomials for their planar mixed link diagrams. Our work provides new tools for the study of annular and toroidal pseudo links.
@@ 0. Introduction
A pseudo knot or link diagram is defined as a planar knot or link diagram containing also double points missing over/under information, called precrossings. These diagrams are classified under generalized Reidemeister moves (see Figure 2) that take into account the indeterminate nature of the precrossings. The equivalence class of a pseudo link diagram is called a pseudo link (see [9, 5] for more details). As presented in our previous work [5], planar, annular and toroidal pseudo link diagrams can lift to spatial pseudo links in the three-sphere, pseudo links in the solid torus, and pseudo links in the thickened torus, respectively. This allows us to translate equivalence of pseudo links diagrams into isotopy of their corresponding lifts. It also allows us to represent pseudo links in the solid torus and in the thickened torus by mixed links in the three-sphere.
@@
Here, we first introduce the notion of regular isotopy for planar, annular and toroidal pseudo link diagrams. Regular isotopy in these diagrammatic settings extends regular isotopy of classical knot and link diagrams, as the invariance under planar isotopy moves, R0, and the Reidemeister moves R2 and R3 [12], by incorporating the additional pseudo Reidemeister moves PR2 and PR3 (see Figure 2). We then define the planar, annular, reduced toroidal and toroidal pseudo bracket polynomials, as $n$-variable Laurent polynomials, with $n=3$ (generalizing the one of [7]), $n=4$ (generalizing the one of [2]), $n=5$ and $n=\infty$, respectively (Definitions 1.42.3 and 3.4), which we prove to be invariants of regular isotopy for the corresponding categories of pseudo links in Propositions 1.6 and  2.5, and Theorem 3.6. By further normalizing the regular isotopy under the Reidemeister move R1 and the pseudo Reidemeister move PR1, we obtain the corresponding Jones-type invariants for planar (recovering the one in [7]), annular and toroidal pseudo links, as well as their three-dimensional lifts in the three-sphere, the solid torus, and the thickened torus, as stated in Theorems 1.8,  2.10, and  3.14. We recapitulate our results in the following theorems.
@@
Theorem 1. The annular pseudo bracket polynomial, resp. the normalized annular pseudo bracket polynomial, is a regular isotopy invariant, resp. an isotopy invariant, of annular pseudo links and their corresponding pseudo links in the solid torus.
@@
Theorem 2. The (reduced) toroidal pseudo bracket polynomial, resp. the normalized (reduced) toroidal pseudo bracket polynomial, is a regular isotopy invariant, resp. an isotopy invariant, of toroidal pseudo links and their corresponding pseudo links in the thickened torus.
@@
Furthermore, the definitions of pseudo links in the solid and in the thickened torus, as lifts of annular and toroidal pseudo links, also allow us to translate the theories of annular and toroidal pseudo links into the theories of ${\rm O}$-mixed pseudo links and ${\rm H}$-mixed pseudo links in the three-sphere $S^{3}$, respectively [5]. More precisely, an ${\rm O}$-mixed pseudo link is a spatial pseudo link in $S^{3}$ that contains a point-wise fixed unknotted component representing the complementary solid torus (see also [2]). Further, an ${\rm H}$-mixed pseudo link is a spatial pseudo link in $S^{3}$ that contains a point-wise fixed Hopf link as a sublink, whose complement is a thickened torus. On the diagrammatic level, these imply that annular and toroidal pseudo links can be viewed as planar ${\rm O}$-mixed and ${\rm H}$-mixed pseudo link diagrams.
@@
We thus introduce the 4-variable mixed bracket polynomial for planar ${\rm O}$-mixed pseudo links (Definition 2.7), which we prove to be an invariant under regular isotopy of ${\rm O}$-mixed pseudo links, equivalent to the annular pseudo bracket polynomial (Theorem 2.8). Furthermore, we introduce the infinite variable mixed bracket polynomial for planar ${\rm H}$-mixed pseudo links (Definition 3.9), and a reduced 5-variable version (Definition 3.11), which we prove to be invariants under regular isotopy of ${\rm H}$-mixed pseudo links, equivalent to the toroidal and the reduced toroidal pseudo bracket polynomials (Theorem 3.12). Finally, in analogy to the annular and toroidal pseudo links, we normalize the mixed pseudo link polynomials under the moves R1 and PR1, and thus obtain Jons-type invariants for annular and toroidal pseudo links through the theory of mixed links (Theorems  2.10 and  3.14).
@@ 1.2. The classical bracket polynomial
[…]
@@
    $\langle$ [Uncaptioned image] $\rangle$ = $A$langle [Uncaptioned image] $\rangle$ $+$ $A^{-1}$langle [Uncaptioned image] $\rangle$, for diagrams that differ locally around a single crossing.
The bracket polynomial is an invariant of regular isotopy for classical (unoriented) links, meaning that it remains unchanged under equivalence induced by the R0, R2 and R3 moves. To obtain an ambient isotopy invariant for classical knots and links, one needs to normalize the bracket polynomial so as to also respect move R1. This normalized version of the bracket polynomial is equivalent to the Jones polynomial of a link. The normalization process involves assigning an orientation to each component of a link $L$, which in turn assigns $+1$ or $-1$ to each crossing of a given diagram, say $D$, of $L$, as shown in the convention in Figure 4.
@@ 1.3. The planar pseudo bracket polynomial
[…]
@@ 1.4. The planar pseudo Jones polynomial
As in the case of the classical bracket polynomial [12] for links in $S^{3}$, one can normalize the pseudo bracket polynomial in order to satisfy the moves R1 and PR1. For the invariance under the move PR1, the specialization $H=1-Vd$ must be forced, as shown in [7]. For the invariance under the move R1, we consider the product of $\langle L\rangle$ by the factor $\left(-A^{-3}\right)^{w(L)}$ as in the normalization of the classical bracket polynomial. Note that we retain the definition of writhe $w(L)$ of a pseudo link diagram $L$, as defined in [7], where precrossings do not contribute to the counting of signs. So the writhe $w(L)$ is defined as in Equation 1.1 and it is an invariant of regular isotopy for pseudo links. Hence we have the following (see [7, Corollary 2]):
@@
where $w(L):=\underset{c\in C(L)}{\sum}\,sgn(c)$, $C(L)$ the set of classical crossings of $L$ and $\langle L\rangle$ denotes the pseudo bracket polynomial of $L$ for $H=1-Vd$, and it is an isotopy invariant of $L$.
@@ 2.1. Preliminaries
Annular pseudo links can be studied in different topological contexts. Recall first that the lift of an annular pseudo link diagram in the thickened annulus $\mathcal{A}\times I$, where $I$ denotes the unit interval $[0,1]$, is defined as a collection of closed curve(s) constrained to the interior of the solid torus ST - which is homeomorphic to $\mathcal{A}\times I$ consisting of embedded discs from which emanate embedded arcs [5]. More specifically, each classical crossing of the diagram is embedded in a sufficiently small 3-ball that lies entirely within $\mathcal{A}\times I$, while precrossings are supported by sufficiently small rigid discs, which are embedded in $\mathcal{A}\times I$. Note also that the simple arcs connecting the (pre)crossings can be replaced by isotopic ones in $\mathcal{A}\times I$. The resulting lift is called a pseudo link in the solid torus. For an illustration of the lift of an annular pseudo link diagram to a pseudo link in the solid torus, we refer to Figure 7.
@@ 2.2. The annular pseudo bracket polynomial
(a) If we have an annular pseudo diagram that contains no essential closed curves, by using the inclusion of a disc in $\mathcal{A}$ (see left-hand illustration in Figure 10), such annular diagrams can be viewed as planar state diagrams and resolve into states as in the planar case.
(b) In contrast, using the inclusion of $\mathcal{A}$ in a disc (see right-hand illustration in Figure 10), essential unknots become null homotopic, so substituting $s^{k}$ by $d^{k-1}=(-A^{2}-A^{-2})^{k-1}$, the annular pseudo bracket polynomial specializes to the planar pseudo bracket of Definition 1.4 for planar pseudo links.
(c) The main difference of our annular pseudo bracket polynomial from the one defined in [2] is that the variable $H$ is specialized to $H=1-Vd$ in [2].
Figure 10. Inclusion relations (cf. [5]): a disc in the annulus and the annulus in a disc.
@@
The annular pseudo bracket polynomial is invariant under regular isotopy of oriented annular pseudo links.
@@
We refer to [2, Proposition 5.5] and [3, Theorem 8] for details on the invariance under the Reidemeister moves R2, R3, and pseudo Reidemeister moves PR2, PR3, noting that in all proofs there was no need of specializing the variable $H$. $\Box$
@@ 2.3. The ${\rm O}$-mixed pseudo bracket polynomial
In this subsection, we consider the planar ${\rm O}$-mixed pseudo link approach to annular pseudo links (recall Subsection 2.1). In view of Theorem 2.2, the annular pseudo bracket polynomial can be adapted to the setting of (oriented) ${\rm O}$-mixed pseudo links. Here the mixed crossings are not subjected to the inductive rules, since ${\rm O}$ must remain fixed throughout. Further, an essential closed curve embedded in the annulus is represented by a mixed Hopf link, as illustrated in Figure 12.
Figure 12. The transition of an essential unknotted curve in $\mathcal{A}$ to an essential unknotted curve in ST, to an ${\rm O}$-mixed pseudo link (cf. [5]).
More precisely we define:
@@
The ${\rm O}$-mixed pseudo bracket is invariant under regular isotopy of (oriented) ${\rm O}$-mixed pseudo links and it is equivalent to the annular pseudo bracket polynomial for annular pseudo links.
@@ 2.4. Jones-type analogues
where $\langle L\rangle_{\mathcal{A}}$ denotes the annular pseudo bracket polynomial of $L$ for $H=1-Vd$, and it is an isotopy invariant of annular pseudo links and their corresponding pseudo links in the solid torus.
@@
where $\langle{\rm O}\cup L\rangle$ denotes the ${\rm O}$-mixed pseudo bracket polynomial of ${\rm O}\cup L$ for $H=1-Vd$, and it is an isotopy invariant of ${\rm O}$-mixed pseudo links and their corresponding spatial ${\rm O}$-mixed pseudo links.
@@
As in the case of the planar pseudo bracket polynomial (Theorem 1.8), multiplying the annular pseudo bracket polynomial with the writhe correction factor $\left(-A^{-3}\right)^{w(L)}$ ensures invariance under Reidemeister moves R1, while retaining invariance under regular isotopy, since the writhe is invariant of regular isotopy. For the invariance under the move PR1, the specialization $H=1-Vd$ must be forced, as argued in [7] and discussed in Subsection 1.4. So, the polynomial $P$ is an invariant of annular pseudo links, analogous to the normalized pseudo bracket for planar pseudo links.
@@
(a) As in the case of planar pseudo links, the variable $V$ witnesses the presence of pseudo crossings. In the absence of precrossings, the polynomials $P_{L}(A,V,s)$ and $P_{{\rm O}\cup L}(A,V,s)$ specialize to the analogue of the normalized bracket (or the Jones polynomial for $A=t^{-1/4}$) for links in the solid torus (cf. for example [11, 14, 4]).

### v2

@@ Abstract.
Pseudo links are equivalence classes under Reidemeister-type moves of link diagrams containing crossings with undefined over and under information. In this paper, we extend the Kauffman bracket and Jones-type polynomials from planar pseudo links to annular and toroidal pseudo links and their respective lifts from the three-space to the solid torus and the thickened torus. Moreover, since annular and toroidal pseudo links can be represented as mixed links in the three-sphere, we also introduce the respective Kauffman bracket and Jones-type polynomials for their planar mixed link diagrams. Our work provides new tools for the study of annular and toroidal pseudo links.
@@ 0. Introduction
A pseudo knot or link diagram is defined as a planar knot or link diagram containing also double points missing over/under information, called precrossings. These diagrams are classified under generalized Reidemeister moves (see Figure 2) that take into account the indeterminate nature of the precrossings. The equivalence class of a pseudo link diagram is called a pseudo link (see [7, 12] for more details). As presented in our previous work [7], planar, annular and toroidal pseudo link diagrams can lift to spatial pseudo links in the three-sphere or three-space, pseudo links in the solid torus, and pseudo links in the thickened torus, respectively. This allows us to translate equivalence of pseudo links diagrams into isotopy of their corresponding lifts. It also allows us to represent pseudo links in the solid torus and in the thickened torus by spatial pseudo mixed links. Spatial pseudo (mixed) links can be viewed as ‘knotted’ embedded graphs in three-space, where the four-valent vertices may be interpreted as precrossings, and whose equivalence and invariants were studied in [17].
@@
Here, we first introduce the notion of regular isotopy for planar, annular and toroidal pseudo link diagrams. Regular isotopy in these diagrammatic settings extends regular isotopy of classical knot and link diagrams, as the invariance under planar isotopy moves, R0, and the Reidemeister moves R2 and R3 [18], by incorporating the additional pseudo Reidemeister moves PR2 and PR3 (see Figure 2). We then define the planar, (universal) annular, (universal, reduced) toroidal pseudo bracket polynomials, as $n$-variable Laurent polynomials, with $n=3$ (generalizing the one of [9]), $n=4$ (generalizing the one of [4]), $n=5$ and $n=\infty$, respectively (Definitions 1.42.3 and 3.4), which we prove to be invariants of regular isotopy for the corresponding categories of pseudo links in Propositions 1.6 and  2.5, and Theorem 3.7. By further normalizing the regular isotopy under the Reidemeister move R1 and the pseudo Reidemeister move PR1, we obtain the corresponding Jones-type invariants for planar (recovering the one in [9]), annular and toroidal pseudo links, as well as their three-dimensional lifts in the three-sphere, the solid torus, and the thickened torus, as stated in Theorems 1.8,  2.10, and  3.16. We recapitulate our results in the following theorems.
@@
Theorem 1. The (universal) annular pseudo bracket polynomial, resp. the normalized annular pseudo bracket polynomial, is a regular isotopy invariant, resp. an isotopy invariant, of annular pseudo links and their corresponding pseudo links in the solid torus.
@@
Theorem 2. The (universal, reduced) toroidal pseudo bracket polynomial, resp. the normalized (universal, reduced) toroidal pseudo bracket polynomial, is a regular isotopy invariant, resp. an isotopy invariant, of toroidal pseudo links and their corresponding pseudo links in the thickened torus.
@@
Furthermore, the definitions of pseudo links in the solid and in the thickened torus, as lifts of annular and toroidal pseudo links, also allow us to translate the theories of annular and toroidal pseudo links into the theories of $\rm O$}-mixed pseudo links and $\rm H$}-mixed pseudo links in the three-sphere $S^{3}$, respectively, cf. [7]. More precisely, an $\rm O$}-mixed pseudo link is a spatial pseudo link in $S^{3}$ that contains a point-wise fixed unknotted component representing the complementary solid torus (see also [4]). Further, an $\rm H$}-mixed pseudo link is a spatial pseudo link in $S^{3}$ that contains a point-wise fixed Hopf link as a sublink, whose complement is a thickened torus. On the diagrammatic level, these imply that annular and toroidal pseudo links can be viewed as planar $\rm O$}-mixed and $\rm H$}-mixed pseudo link diagrams.
@@
We thus introduce the 4-variable mixed bracket polynomial for planar $\rm O$}-mixed pseudo links (Definition 2.7) and its universal analogue, which we prove to be invariants under regular isotopy of $\rm O$}-mixed pseudo links, equivalent to the (universal) annular pseudo bracket polynomial (Theorem 2.8). Furthermore, we introduce the infinite variable (universal) mixed bracket polynomial for planar $\rm H$}-mixed pseudo links (Definition 3.11), and a reduced 5-variable version (Definition 3.13), which we prove to be invariants under regular isotopy of $\rm H$}-mixed pseudo links, equivalent to the (universal) toroidal and the reduced toroidal pseudo bracket polynomials (Theorem 3.14). Finally, in analogy to the annular and toroidal pseudo links, we normalize the (universal) mixed pseudo link polynomials under the moves R1 and PR1, and thus obtain Jones-type invariants for annular and toroidal pseudo links through the theory of mixed links (Theorems  2.10 and  3.16).
@@ 1.2. The classical bracket polynomial
The bracket polynomial, denoted by $\langle\cdot\rangle$, was introduced by Kauffman in [18] as a one-variable Laurent polynomial in the ring $\mathbb{Z}[A^{\pm 1}]$ defined recursively by means of the following rules:
@@
    $\langle$ [Uncaptioned image] $\rangle$ = AA $\langle$ [Uncaptioned image] $\rangle$ ++ $A^{-1}$ $\langle$ [Uncaptioned image] $\rangle$, for diagrams that differ locally around a single crossing.
The bracket polynomial is an invariant of regular isotopy for classical (unoriented) links, meaning that it remains unchanged under equivalence induced by the R0, R2 and R3 moves. To obtain an ambient isotopy invariant for classical knots and links, one needs to normalize the bracket polynomial so as to also respect move R1. This normalized version of the bracket polynomial is equivalent to the Jones polynomial of a link. The normalization process involves assigning an orientation to each component of a link $L$, which in turn assigns +1+1 or 1-1 to each crossing of a given diagram, say $D$, of $L$, as shown in the convention in Figure 4.
@@ 1.3. The planar pseudo bracket polynomial
In [9] the 2-variable pseudo bracket polynomial is defined for planar pseudo link diagrams, extending the Kauffman bracket polynomial $\langle\cdot\rangle$ for classical knots and links [18]. This extension was achieved by considering oriented pseudo link diagrams, which allowed the definition of a skein relation at a precrossing. Note that this skein relation can be viewed as an operation where specific tangles are inserted to replace the precrossings. See [13] for the application of tangle insertion to pseudo knots. In the setting of [9], which aimed at an ambient isotopy invariant, the pseudo bracket polynomial was forced to satisfy the move PR1. However, the move PR1 is not included in our definition of regular isotopy for pseudo links (Definition 1.1), so we define the pseudo bracket polynomial as a 3-variable polynomial as follows:
@@ 1.4. The planar pseudo Jones polynomial
As in the case of the classical bracket polynomial [18] for links in $S^{3}$, one can normalize the pseudo bracket polynomial in order to satisfy the moves R1 and PR1. For the invariance under the move PR1, the specialization H=1VdH=1-Vd must be forced, as shown in [9]. For the invariance under the move R1, we consider the product of $\langle L\rangle$ by the factor $\left(-A^{-3}\right)^{w(L)}$ as in the normalization of the classical bracket polynomial. Note that we retain the definition of writhe $w(L)$ of a pseudo link diagram $L$, as defined in [9], where precrossings do not contribute to the counting of signs. So the writhe $w(L)$ is defined as in Equation 1.1 and it is an invariant of regular isotopy for pseudo links. Hence we have the following (see [9, Corollary 2]):
@@
where $w(L):=\underset{c\in C(L)}{\sum}\,sgn(c)$, $C(L)$ the set of classical crossings of $L$ and $\langle L\rangle$ denotes the pseudo bracket polynomial of $L$ for H=1VdH=1-Vd, and it is an isotopy invariant of $L$.
@@ 2.1. Preliminaries
Annular pseudo links can be studied in different topological contexts. Recall first that the lift of an annular pseudo link diagram in the thickened annulus $\mathcal{A}\times I$, where $I$ denotes the unit interval $[0,1]$, is defined as a collection of closed curve(s) constrained to the interior of the solid torus ST - which is homeomorphic to $\mathcal{A}\times I$ - consisting of embedded discs from which emanate embedded arcs [7]. More specifically, each classical crossing of the diagram is embedded in a sufficiently small 3-ball that lies entirely within $\mathcal{A}\times I$, while precrossings are supported by sufficiently small rigid discs, which are embedded in $\mathcal{A}\times I$. Note also that the simple arcs connecting the (pre)crossings can be replaced by isotopic ones in $\mathcal{A}\times I$. The resulting lift is called a pseudo link in the solid torus. For an illustration of the lift of an annular pseudo link diagram to a pseudo link in the solid torus, we refer to Figure 7.
@@ 2.2. The annular pseudo bracket polynomial
(a) If we have an annular pseduo diagram that contains no essential closed curves, by using the inclusion of a disc in $\mathcal{A}$ (see left-hand illustration in Figure 10), such annular diagrams can be viewed as planar state diagrams and resolve into states as in the planar case.
(b) In contrast, using the inclusion of $\mathcal{A}$ in a disc (see right-hand illustration in Figure 10, cf. [7]), essential unknots become null homotopic, so substituting $s^{k}$ by $d^{k-1}=(-A^{2}-A^{-2})^{k-1}$, the annular pseudo bracket polynomial specializes to the planar pseudo bracket of Definition 1.4 for planar pseudo links.
(c) The main difference of our annular pseudo bracket polynomial from the one defined in [4] is that the variable $H$ is specialized to H=1VdH=1-Vd in [4].
(d) In the fifth rule of Definition 2.3 the initial condition could be substituted by $s_{k}$ in place of $s^{k}$, for $k\in\mathbb{N}$, giving rise to the universal annular pseudo bracket polynomial with infinitely many variables. Since precrossings are also smoothened, the universal annular pseudo bracket polynomial recovers the Kauffman bracket skein module of annular pseudo links or pseudo links in the solid torus, cf. [15, 4].
Figure 10. Inclusion relations: a disc in the annulus and the annulus in a disc.
@@
The (universal) annular pseudo bracket polynomial is invariant under regular isotopy of oriented annular pseudo links.
@@
We refer to [4, Proposition 5.5] for details on the invariance under the Reidemeister moves R2, R3, and pseudo Reidemeister moves PR2, PR3, noting that in all proofs there was no need of specializing the variable $H$. $\Box$
@@ 2.3. The $\rm O$}-mixed pseudo bracket polynomial
In this subsection, we consider the planar $\rm O$}-mixed pseudo link approach to annular pseudo links (recall Subsection 2.1). In view of Theorem 2.2, the annular pseudo bracket polynomial can be adapted to the setting of (oriented) $\rm O$}-mixed pseudo links. Here the mixed crossings are not subjected to the inductive rules of Definition 2.3, since $\rm O$} represents the manifold, so must remain fixed throughout. Further, an essential closed curve embedded in the annulus is represented by a mixed Hopf link, as illustrated in Figure 12 (cf. [19], see also [7]).
Figure 12. The transition of an essential unknotted curve in $\mathcal{A}$ to an essential unknotted curve in the solid torus, to an $\rm O$}-mixed pseudo link.
In the case of $\rm O$}-mixed pseudo links it may look as though a generic state diagram includes finitely many essential unknotted components winding around the fixed $\rm O$}-curve some number of times, as illustrated in Figure 13. However, in the Kauffman bracket skein module of the solid torus such diagrams reduce to diagrams with essential unknotted components winding once around the fixed $\rm O$}-curve, as illustrated in Figure 14. Cf. [6, 10].
Figure 13. A state diagram with $k_{1},\ldots,k_{n}$ identical essential unknots, winding $a_{1},\ldots,a_{n}$ times around the $\rm O$}-curve, respectively.
We may now define:
@@
Further, in analogy to the universal annular pseudo bracket polynomial we have the universal $\rm O$}-mixed pseudo bracket polynomial with infinitely many variables $s_{k}$, for $k\in\mathbb{N}$, in place of $s$.
@@
The (universal) $\rm O$}-mixed pseudo bracket is invariant under regular isotopy of (oriented) $\rm O$}-mixed pseudo links and it is equivalent to the (universal) annular pseudo bracket polynomial for annular pseudo links.
@@ 2.4. Jones-type analogues
where $\langle L\rangle_{\mathcal{A}}$ denotes the annular pseudo bracket polynomial of $L$ for H=1VdH=1-Vd, and it is an isotopy invariant of annular pseudo links and their corresponding pseudo links in the solid torus. Likewise for the universal normalized annular pseudo bracket polynomial $P_{{\rm O}\cup L}(A,V,s_{k})$ for $k\in\mathbb{N}$.
@@
where $\langle{\rm O}\cup L\rangle$ denotes the $\rm O$}-mixed pseudo bracket polynomial of ${\rm O}\cup L$ for H=1VdH=1-Vd, and it is an isotopy invariant of $\rm O$}-mixed pseudo links and their corresponding spatial $\rm O$}-mixed pseudo links. Likewise for the universal normalized $\rm O$}-mixed pseudo bracket polynomial $P_{{\rm O}\cup L}(A,V,s_{k})$ for $k\in\mathbb{N}$.
@@
As in the case of the planar pseudo bracket polynomial (Theorem 1.8), multiplying the annular pseudo bracket polynomial with the writhe correction factor $\left(-A^{-3}\right)^{w(L)}$ ensures invariance under Reidemeister moves R1, while retaining invariance under regular isotopy, since the writhe is invariant of regular isotopy. For the invariance under the move PR1, the specialization H=1VdH=1-Vd must be forced, as argued in [9] and discussed in Subsection 1.4. So, the polynomial $P$ is an invariant of annular pseudo links, analogous to the normalized pseudo bracket for planar pseudo links.
@@
(a) As in the case of planar pseudo links, the variable $V$ witnesses the presence of precrossings. In the absence of precrossings, the polynomials $P_{L}(A,V,s)$ and $P_{{\rm O}\cup L}(A,V,s)$ specialize to the analogue of the normalized bracket (or the Jones polynomial for $A=t^{-1/4}$) for links in the solid torus (cf. [15, 19]).